OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (anonymous):

\[((sinx)/(1-cosx)) + ((sinx)/(1+cosx)) = 2 \csc x\]

zepdrix (zepdrix):

Ooo these are fun. Hmmm let's see.. You remember how to do stuff with cojugates? (a-b)(a+b)

zepdrix (zepdrix):

If we multiply the first fraction, top and bottom, by the `conjugate` of the denominator, and then do the same with the other fraction, we can get a nice common denominator.

zepdrix (zepdrix):

\(\large\rm \dfrac{\sin x}{1-\cos x}\color{royalblue}{\left(\dfrac{1+\cos x}{1+\cos x}\right)}+\dfrac{\sin x}{1+\cos x}\color{orangered}{\left(\dfrac{1-\cos x}{1-\cos x}\right)}\)

zepdrix (zepdrix):

Don't multiply out the tops, leave the brackets. Multiplying out the bottoms gives us the difference of squares \(\Large\rm 1-\cos^2x\). Do you remember your Pythagorean Identity for Sine and Cosine?

OpenStudy (anonymous):

sin=o/ad cos=a/h @zepdrix

zepdrix (zepdrix):

No no no you silly billy :o Our Pythagorean Identity is: \(\Large\rm \sin^2x+\cos^2x=1\) If we subtract cos^2x from each side it gives us, \(\Large\rm \sin^2x=1-\cos^2x\) Do you see how we can maybe use this? :o

OpenStudy (anonymous):

oh yea oops wrong one :/

zepdrix (zepdrix):

After multiplying out the conjugates our equation is currently, \(\Large\rm \dfrac{\sin x(1+\cos x)}{1-\cos^2x}+\dfrac{\sin x(1-\cos x)}{1-\cos^2x}\)

OpenStudy (anonymous):

Would I somehow combine them?

zepdrix (zepdrix):

Not just yet. The Pythagorean Identity allows us to convert our denominators to sin^2x, yes?

OpenStudy (anonymous):

yes since sin^2x=1−cos^2x

zepdrix (zepdrix):

After that you can cancel out a sin from each top and bottom, leaving you with: \(\Large\rm \dfrac{1+\cos x}{\sin x}+\dfrac{1-\cos x}{\sin x}\)And from THIS place, it would be a good idea to combine.

zepdrix (zepdrix):

\(\Large\rm \dfrac{1+\cos x+1-\cos x}{\sin x}\)

zepdrix (zepdrix):

What should we do next? :o What do you think?

OpenStudy (anonymous):

2/sinx?

zepdrix (zepdrix):

Ok good. Let's pull the 2 off of the fraction so maybe you'll remember another identity we can use. \(\Large\rm 2\left(\dfrac{1}{\sin x}\right)\)

zepdrix (zepdrix):

Bah I've gotta get going :c